# OFDM-Based Active STAR-RIS-Aided Integrated Sensing and Communication Systems

### v1

@@ VI-E The effect of different number of OFDM subcarriers in the velocity estimation
VI-F QPSK BER and high-velocity circumstances
In Fig. $9$, we compare the quadrature phase-shift keying (QPSK) BER across different true velocities. It is clear that as $v_{\textup{true}}$ increases, the BER also increases. This indicates that the data estimation performance deteriorates under high-velocity conditions. Consequently, in Fig. $10$, we adhere to the 5G NR criteria, observing that an increase in subcarrier spacing results in a decrease in BER. This suggests that the data estimation performance in high-speed conditions can be improved by increasing the subcarrier spacing. Increasing the subcarrier spacing reduces the impact of Doppler shifts, which are more pronounced at higher velocities. This mitigation of Doppler effects is crucial for maintaining the integrity of signal transmission, also enhancing data estimation performance in high-speed scenarios.
Note that our BER appears to follow a linear scale; this is because the BER ranges within a narrow interval, making it challenging to represent on a logarithmic scale. Despite the limited BER range, it still effectively demonstrates the desired properties we aim to investigate. This phenomenon has also been discussed in previous works, such as [1, 37] and [38].
VII Conclusion
@@
$\geq\xi\big{(}|\mathbf{u}_{k}^{H}(n)\mathbf{W}_{r}|_{2}^{2}+\sum_{i\neq k}\mathbf{u}_{k}^{H}(n)\mathbf{R}_{i}\mathbf{u}_{k}$
@@
From (32), we use the equation $A(n)=|\mathbf{u}_{k}^{H}(n)\mathbf{W}_{r}|_{2}^{2}+\sum_{i\neq k}\mathbf{u}_{k}^{H}(n)\mathbf{R}_{i}\mathbf{u}_{k}+\sigma_{v}^{2}\mathbf{t}_{k}^{H}(n)\boldsymbol{\Psi}_{t}\boldsymbol{\Psi}_{t}^{H}\mathbf{t}_{k}(n)+\sigma^{2}$ to simpify the term on the right side

### v3

@@ V-E The effect of different number of OFDM subcarriers in the velocity estimation
V-F BPSK BER and high-velocity circumstances
In Fig. $9$, we compare the binary phase-shift keying (BPSK) BER due to inter-carrier interference (ICI) across different true velocities. The generation of ICI in our simulation is achieved by introducing phase and amplitude errors that scale with the Doppler frequency shift. Specifically, the phase error is modeled as a random variable uniformly distributed between $-\pi$ and $\pi$, scaled by an ICI effect factor. The amplitude error is modeled as a small perturbation scaled by the square root of the ICI effect factor. These errors are applied to each subcarrier, simulating the ICI experienced in real-world high-speed scenarios. It is clear that as $v_{\textup{true}}$ increases, the BER also increases. This indicates that the data estimation performance deteriorates under high-velocity conditions. Consequently, in Fig. $10$, we adhere to the 5G NR criteria, observing that an increase in subcarrier spacing results in a decrease in BER. This suggests that the data estimation performance in high-speed conditions can be improved by increasing the subcarrier spacing. Increasing the subcarrier spacing reduces the impact of Doppler shifts, which are more pronounced at higher velocities. This mitigation of Doppler effects is crucial for maintaining the integrity of signal transmission, also enhancing data estimation performance in high-speed scenarios. In Fig. $11$, we have plotted the number of RIS elements versus BPSK BER. The results show that when we increase the number of RIS elements, the BPSK BER decreases simultaneously. This occurs because more RIS elements enhance the signal reflection and focusing capabilities, leading to a stronger and more coherent signal at the receiver. Consequently, the improved signal quality reduces the BER, resulting in more reliable communication. Additionally, we can confirm that when the subcarrier spacing increases, the BER decreases at the same time, even when the horizontal axis represents the number of RIS elements. This is because larger subcarrier spacing reduces inter-carrier interference, further improving signal clarity. These combined effects demonstrate the importance of optimizing both RIS configuration and subcarrier spacing for minimizing BER.
VI Conclusion
@@
$\geq\xi\big{(}|\mathbf{u}_{k}^{H}(n)\mathbf{W}_{r}|_{2}^{2}+\sum_{i\neq k}\mathbf{u}_{k}^{H}(n)\mathbf{R}_{i}\mathbf{u}_{k}(n)$
@@
From (33), we use the equation $A(n)=|\mathbf{u}_{k}^{H}(n)\mathbf{W}_{r}|_{2}^{2}+\sum_{i\neq k}\mathbf{u}_{k}^{H}(n)\mathbf{R}_{i}\mathbf{u}_{k}(n)+\sigma_{v}^{2}\mathbf{t}_{k}^{H}(n)\boldsymbol{\Psi}_{t}\boldsymbol{\Psi}_{t}^{H}\mathbf{t}_{k}(n)+\sigma^{2}$ to simpify the term on the right side